OpenStudy (anonymous):

[8.03] Factor completely: x2 − 4x + 9 (x − 1)(x − 9) (x − 3)(x − 3) (x + 3)(x − 3) Prime I think its Prime help me please

OpenStudy (nurali):

x²-4x+9 = (x²-4x+4) +5 = (x-2)²-4+5 = (x-2)² +1 I think its Prime

OpenStudy (anonymous):

that's what i think

OpenStudy (nurali):

Prime

OpenStudy (anonymous):

ok

OpenStudy (nurali):

if correct please click best response.